# CAIN: Hijacking LLM-Humans Conversations via a Two-Stage Malicious System Prompt Generation and Refining Framework

### v1

@@ 5.1 Setup
Metrics. We use two sets of metrics, including (1) Predictive F1 and Exact Match (EM): standard Q&A metrics measuring partial and exact correctness of model prediction against ground-truths, and (2) Performance gap $\Delta$F1 and $\Delta$EM measure the difference in performance between benign and malicious tasks (e.g., $\Delta\text{F1}{=}\text{F1}_{benign}{-}\text{F1}_{malicious}$). A higher $\Delta$F1/EM indicates a stronger attack, meaning a greater performance drop on the target set with minimal loss on the benign set.
Target LLMs and Attack Baselines. We evaluate on six open-source LLMs, including Llama2 Touvron et al. (2023), LLama3.1 Grattafiori et al. (2024), Deepseek DeepSeek-AI et al. (2025), Qwen Qwen et al. (2025), Pythia Biderman et al. (2023) with the black-box attack baselines:
  • No system prompt (NSP): Questions are fed to LLMs without any instructions.
  • Manual: A hand-crafted Q&A system prompt.
  • AdvAutoPrompt (AAP): Partially malicious prompt produced by a customized adversarial version of AutoPrompt Levi et al. (2024) formulated in Sec. 4.1.
  • CAIN: Our proposed attack method that combines AAP with greedy word-level optimization.
5.2 Results
@@
Table 1: Performance comparison when attacking various target LLMs. Bold and underlined values indicate the best and second-best $\Delta$F1/EM improvements, respectively.
@@
Table 2: Performance on the targeted attack across various target LLMs. Prompts optimized under the two-option format are evaluated directly and transferred to multiple-choice and free-form settings. Bold indicates best, underline indicates second-best per block.
Figure 3: Ablation study on varying the threshold $k$ and numbers of benign questions on Llama2-13B.
Table 1 reports F1 and EM on Benign and Malicious Evaluation sets. Key findings include: (1) CAIN consistently demonstrates superior adversarial performance on malicious tasks across models, with notably low F1 and EM scores, even with paraphrased versions of the target question, (2) AAP exhibits strong malicious F1 compared to Manual on most of target LLMs, although increased malicious scores on Llama2-7B, and (3) NSP and Manual prompts show strong robustness on benign and malicious task with high F1 scores.
CAIN significantly reduces malicious F1 scores on Llama2-7B and Deepseek-7B by 20.64 and 71.85, respectively, while only modestly affecting benign F1 (drops of 9.24 and 8.12), compared to Manual (Table 1). For Llama2-13B and Qwen2.5-7B, benign performance drops by  20 points, possibly due to the limited number of benign samples. However, Figure 3 (right) shows that increasing benign questions does not improve performance, which we leave for future study.
6 Experiments: Targeted Attack
6.1 Setup
Dataset and Data Sampling (by the Attackers): We use six categories from TruthfulQA (Lin et al., 2022): Misconceptions, Conspiracy, Stereotype, Health, Politics, History. The statistical information is shown in Table A1 (Appendix). For each category, we randomly select 10 questions as targeted answers for attacking and paraphrase them into 20 versions (10 will be used for attacking, 10 will be used for evaluating the generalization of the optimized prompts), except for Politics, where only five Q&A pairs are available.
@@ 6.1 Setup
  • Benign Evaluation: For each category, 5 other questions are selected and paraphrased into 50 variants to evaluate generalization on unseen but non-targeted queries.
  • Malicious Evaluation: The other 10 paraphrased versions of each target question as mentioned above are used for evaluation.
@@
Target LLMs and Attack Baselines. We utilize the same attack baselines as in the untargeted attack setting (§$\S$5). All experiments are conducted on Deepseek-7B, Qwen2.5-7B, and Llama2-7B. [^1] Due to space, Llama2-7B’s results are in the Appendix
Evaluation Settings. Models are prompted to select one answer from different answering formats of increasing difficulties: (1) Two options (A or B), (2) Four options (A, B, C or D), and (3) Free-form text (no explicit choices are provided). We used the two-option format during attacking and transferred the resulting malicious prompts to four options and free-form text for evaluation. Figure A1 illustrates the input format for two options.
6.2 Results
We report results for the two-option setting, its transfer to four-option, and to free-form generation in Table 2. Key findings include: (1) combining AdvAutoPrompt with greedy word-level optimization consistently achieves superior overall performance in benign and malicious responses, highlighting its strong adversarial power, and (2) strong resilience and adaptability to different levels of difficulty, with greedy decoding consistently achieved high performance across evaluation scenarios, emphasizing the crucial role of this optimization strategy.
@@ 7 Discussion
Affects of Optimization Threshold $k$. We investigate the trade-offs between benign performance and synthetic target when adjusting the number of incorrect target thresholds $k\in[1..11]$ (Alg. 1, Ln.. 17). As shown in Figure 4, increasing $k$ consistently improves attack effectiveness while benign performance remains stable. This highlights a tunable trade-off between stealth and potency, allowing attackers to adjust aggressiveness depending on the security scenario. The full analysis is in A.3.
Affects of Model Sizes.
Figure 5: Performance of each attacking scenario across model sizes.
We evaluate how model size impacts CAIN’s effectiveness using Qwen2.5 with the number of parameters increasing from 3B to 32B. As shown in Figure 5 and 3 (left), CAIN consistently achieves stronger adversarial performance than AAP across both targeted and untargeted attacks, with benign performance improving as model size increases. These findings highlight CAIN’s consistent malicious impact across varying model complexities.
Affects of Prompt Initialization Methods. Across both untargeted and targeted settings, A+Greedy consistently outperforms M+Greedy in attack effectiveness and robustness. In untargeted attacks (Table A3), A+Greedy yields stronger performance gaps (e.g., +30.48 $\Delta$F1 on L2-7B, +15.84 on D-7B), indicating better degradation of malicious responses while preserving benign ones. In targeted attacks (Table A5), it achieves higher aggregated scores ($\Psi$), particularly on Qwen2.5-7B and Deepseek-7B, showing better generalization across formats. These results confirm A+Greedy as a superior initialization strategy for attack strength and transferability.
7.1 Case Study: Attacking Commercial LLMs
@@
4.1-nano Manual 62.47 51.03 95.00 95.00 -32.53 -43.97
AAP 64.43 54.25 97.53 93.94 -33.10 -39.69
CAIN 56.79 42.43 81.76 65.00 -24.97 -22.57
3.5Turbo Manual 69.15 51.52 99.55 99.55 -30.40 -48.03
@@
Table 3: Untargeted performance across different attack methods on various OpenAI APIs.
We evaluate CAIN under untargeted attacks on popular commercial LLM-based chatbot APIs, including GPT3.5-turbo, GPT-4.1-nano and GPT-4.1-mini. Due to budget constraints, we randomly selected 10 target questions to attack. Table 3 shows that CAIN consistently reduces F1 score on malicious tasks by 46.84 points for 4o-mini, compared to the manual prompt, while better preserving benign performance by increasing to 71.44 F1.
Moreover, compared to AAP, CAIN consistently achieves lower malicious F1 across models (e.g., 69.47 vs. 96.57 on GPT-3.5-Turbo; 81.76 vs. 97.53 on GPT-4.1-nano), highlighting stronger attack success. While GPT-4.1-nano appears more robust ($\Delta$F1 of -24.97 for CAIN vs. -33.10 for AAP), the results demonstrate CAIN’s effectiveness even against well-aligned commercial models under black-box conditions, confirming the feasibility of our security threat in practice.
TargetLLM Targeted UnTargeted
Deepseek-7B 0.0217 0.0819
Qwen2.5-7B 0.0426 0.0417
Table 4: Averaged cosine similarity between successfully optimized prompts and the targeted questions.
Potential Defense. Our findings suggest that traditional defenses, such as detecting lexical similarity between prompts and target questions or using a perplexity-based filtering Jain et al. (2023) are insufficient to defend against CAIN. Table 4 shows that the optimized prompts exhibit very low cosine similarity to their respective targets (average of 0.0518 for Deepseek-7B and 0.04215 for Qwen2.5-7B), indicating they do not leak any lexical overlap with the target questions. Figure A2 shows that CAIN’s prompts have variable perplexity levels, and using a PPL filtering might work to some extent, but this approach will not be a comprehensive solution. These results underscore CAIN’s subtlety and highlight the urgent need for more robust, behavior-based detection mechanisms.
8 Conclusion
We introduce CAIN, a black-box method that reveals a new vulnerability in LLMs: targeted prompt-based manipulation that preserves benign behavior. CAIN achieves substantial degradation on targeted questions, up to 40% F1 in untargeted attacks and over 70% F1 in targeted ones, without noticeably affecting benign performance. These attacks remain stealthy, transferable across model architectures, and evade traditional defenses such as lexical similarity or perplexity-based filtering.
Limitation
While CAIN demonstrates strong targeted manipulation in black-box settings, it faces several important limitations. First, achieving high adversarial effectiveness occasionally comes at the cost of benign performance. CAIN outperforms baselines on OpenAI APIs, the overall attack success remains limited due to alignment constraints in commercial systems. Finally, while CAIN evades common lexical and perplexity-based filters, this also underscores a broader limitation in the field: the lack of robust, behavior-aware defenses. Addressing these challenges will be crucial for advancing both offensive and defensive research in LLM alignment.
Broader Impacts and Ethics Statement
This work reveals a previously underexplored vulnerability in large language models (LLMs): the ability to craft adversarial system prompts that selectively cause incorrect responses to specific questions while maintaining accurate outputs on benign inputs. Such selective manipulation poses a subtle but serious threat, particularly in domains involving misinformation, political influence, or public health. Unlike traditional jailbreaks or universal attacks, CAIN operates stealthily, evading detection by standard lexical similarity and perplexity filters.
We intend to raise awareness of this threat and prompt the development of more robust, behavior-based defenses. All experiments were conducted in controlled settings using open-source models, and evaluations on commercial APIs were performed to assess practical limitations — not for misuse. While the techniques may be misused, we believe that exposing this vector responsibly contributes to a more secure and trustworthy deployment of LLMs. We advocate for responsible disclosure, transparent benchmarking, and the implementation of proactive safeguards in future LLM systems.
@@ A.1 Dataset and Prompt template
Table A1 shows the statistical information of each category in the TruthfulQA Lin et al. (2022), which will be used for targeted attack. Figure A1 illustrates an example of the input for targeted attack during training and evaluation.
@@ A.2 Implementation Details
For each targeted question, we used 20 benign examples (10 from the correct subset, 10 from the incorrect subset) and the correct proportion $m=7$. For the AdvAutoPrompt, we set a maximum length of $p^{*}$ up to 150 words and the number of iterations $t=3$ to save cost.
@@ A.3 Analysis of optimization threshold $k$
For untargeted attack (top row), increasing the incorrect targeted threshold ($k$) generally leads to a more substantial degradation of performance on the synthetic target (malicious) dataset for both models. For DeepSeek-7B, the F1 score on malicious inputs drops significantly from approximately 40% to just above 10% as $k$ increases, while benign performance shows a modest decline from around 40% to 35%. Qwen2.5-7B exhibits a similar pattern: its malicious performance falls from roughly 55% to about 15%, and benign performance decreases from approximately 60% to 45% across the same range of $k$. This trend indicates that while the attack more effectively degrades performance on malicious queries at higher k values, this also comes with a slight decrease in performance on benign questions compared to their corresponding initial attacked state and a more noticeable reduction compared to the NSP benign baselines.
Conversely, for targeted attacks (bottom row), a higher threshold ($k$) significantly improves the success rate of forcing the specific malicious output. On DeepSeek-7B, the malicious F1 score, representing attack success, climbs from about 58% to nearly 80% with an increasing $k$, while performance on benign inputs remains relatively stable around 50%, comparable to its NSP benign baseline. A more pronounced trend is observed for Qwen2.5-7B, where its malicious attack success rate rises from approximately 50% at $k=1$ to over 80% for $k\geq 8$; its benign performance also remains stable at around 62%. Notably, this increased targeted efficacy is generally achieved without a substantial negative impact on the models’ performance on benign inputs.
Figure A1: The prompt template for two options of targeted attack.
@@ A.4 Additional results on untargeted attack
Table A2 presents the performance of untargeted attacks on Llama3.1-7B and Pythia-12B. Across both models, CAIN significantly outperforms all baselines, including Manual and AAP, in balancing attack strength and benign performance. While Manual prompts achieve high benign F1/EM, they fail to reduce malicious performance (e.g., 99.75 F1 on Llama3.1 and 100.00 F1 on Pythia). In contrast, CAIN reduces malicious F1 to 27.46 and 32.32, respectively, while maintaining reasonable benign scores. This results in the highest $\Delta$F1 and $\Delta$EM margins (e.g., +17.69 F1 on Llama3.1 and +16.76 on Pythia), demonstrating CAIN’s superior ability to selectively degrade targeted outputs without broadly compromising accuracy.
@@
Table A3: Results of attacking performance with manual initialization and AdvAutoPrompt. “A” denotes AAP, “G” stands for Greedy, and “A+G” is our proposed method. “L, D, Q, P” denote Llama, Deepseek, Qwen, and Pythia models, respectively.
@@ A.5 Additional results for targeted attacks
Table A4 shows that AAP achieves the best performance in the two-option setting ($\Psi$F1 = 50.94, $\Psi$EM = 42.58), but its effectiveness drops when transferred to the four-option format. In contrast, CAIN maintains more stable performance across both settings, achieving strong targeted attack success (highest malicious F1) with better transferability ($\Psi$F1 = 31.01 vs. 32.80). This suggests CAIN is more robust and generalizable under realistic conditions where question formats vary.
@@
Benign Malicious Sum Benign Malicious Sum
@@
Llama2-7B NSP 10.03 1.67 14.09 3.27 12.06 2.47 9.72 1.33 19.60 1.82 14.66 1.58
Manual 19.67 19.00 44.03 43.45 31.85 31.23 1.00 1.00 3.50 3.27 2.25 2.13
M+Greedy 28.20 21.57 60.71 58.00 44.45 39.78 13.73 10.60 29.81 28.80 21.77 19.70
AAP 42.47 34.16 59.41 51.00 50.94 42.58 27.30 19.88 38.29 33.27 32.80 26.58
CAIN 35.10 19.61 61.86 48.71 48.48 34.16 20.44 11.83 41.58 34.84 31.01 23.34
Table A4: Performance of the targeted attack on Llama2-7B.
@@
Benign Malicious Sum ($\Psi$) Benign Malicious Sum ($\Psi$)
@@
Deepseek-7B M+Greedy 47.94 46.11 42.90 42.26 45.42 44.19 27.38 25.65 30.32 29.35 28.85 27.50
A+Greedy 55.29 46.47 58.92 54.00 57.11 50.23 31.73 28.69 43.92 43.00 37.83 35.84
Qwen2.5-7B M+Greedy 60.41 60.33 62.73 62.73 61.57 61.53 50.88 50.88 69.70 69.70 60.29 60.29
A+Greedy 62.76 61.75 73.03 72.07 67.89 66.91 56.45 56.38 70.75 70.34 63.60 63.36
Table A5: Performance with different initialization methods on targeted attacks.

### v2

@@ 5 Untargeted Attack
[…]
Target LLMs and Attack Baselines. We evaluate on six open-source LLMs, including Llama2 (Touvron et al. 2023), LLama3.1 (Dubey, Jauhri, and Authors 2024), Deepseek (DeepSeek-AI 2025), Qwen (Qwen 2025), Pythia (Biderman et al. 2023) with baselines: (1) NSP: no system prompt, (2) Manual: a hand-crafted QA prompt (Table A.4), (3) AdvAutoPrompt (AAP): the partially malicious prompt from Stage 1, and (4) CAIN: our proposed method combining AAP with greedy word-level optimization.
Results. Table 2 reports F1 and EM on Benign and Malicious Evaluation sets. Key findings include: (1) CAIN consistently demonstrates superior adversarial performance on malicious tasks across models, with notably low F1 and EM scores, even with paraphrased versions of the target question, (2) AAP exhibits strong malicious F1 compared to Manual on most of target LLMs, although increased malicious scores on Llama2-7B, and (3) NSP and Manual prompts show strong robustness on benign and malicious task with high F1 scores.
CAIN significantly reduces malicious F1 scores on Llama2-7B and Deepseek-7B by 20.64 and 71.85, respectively, while only modestly affecting benign F1 (drops of 9.24 and 8.12), compared to Manual (Table 2). For Llama2-13B and Qwen2.5-7B, benign performance drops by $\approx 20$ points. This drop is expected because the goal of $\mathcal{Q}_{b}$ is not to improve benign performance, but to maintain a sufficient utility and prevent the adversarial prompt from excessively degrading it. In practice, when we optimize for a specific target question (e.g., “Are COVID vaccines safe?”), we are implicitly optimizing for semantically related questions (e.g., “Are vaccines safe?”), which can shift the benign distribution and influence the model’s optimization direction. Consequently, modifying $\mathcal{Q}_{b}$ changes the distribution of benign queries, which in turn affects how the model adjusts to preserve benign performance along those dimensions.
This intuition is further confirmed by Fig. 4(right), which shows that increasing the size of $\mathcal{Q}_{b}$ does not lead to a better F1 score. We attribute this to an inherent trade-off between targeted degradation and benign preservation, rather than to insufficient data. Simply increasing the size of $\mathcal{Q}_{b}$ may not improve benign utility if the added samples do not contribute meaningful new constraints. Another hypothesis is that benign performance is strongly dependent on specific target questions, as manipulating general questions such as “Are vaccines dangerous?” is different from a more specific question like “Are COVID vaccines dangerous?”. A deeper investigation of this behavior is left for future work.
@@
[…]
Benign Malicious Difference Benign Malicious Difference Benign Malicious Difference
@@
Table 3: Targeted attack performance across target LLMs. Prompts optimized in the two-option format are evaluated directly and transferred to multiple-choice and free-form settings. Bold and underline denote the best and second-best results per block.
Figure 4: Benign (blue) and malicious (red) performance with varying $k$ and numbers of benign questions on Llama2-13B.
6 Targeted Attack
Dataset and Data Sampling (by the Attackers): We use six categories from TruthfulQA (Lin, Hilton, and Evans 2022): Misconceptions, Conspiracy, Stereotype, Health, Politics, History. The statistical information is shown in Table A1 (Appendix). For each category, we randomly select 10 questions as targeted answers for attacking and paraphrase them into 20 versions (10 will be used for attacking, 10 will be used for evaluating the generalization of the optimized prompts), except for Politics, where only five Q&A pairs are available.
@@ 6 Targeted Attack
  • Benign Evaluation: For each category, 5 other non-targeted questions are paraphrased into 50 variants to evaluate generalization on unseen benign queries.
  • Malicious Evaluation: Other 10 paraphrased versions of each target question are used to assess attack effectiveness.
@@
Target LLMs and Attack Baselines. We adopt the same attack baselines as in the untargeted attack setting and evaluated on Deepseek-7B, Qwen2.5-7B, and Llama2-7B. [^1] Due to space, Llama2-7B’s results are in the Appendix
Evaluation Settings. Models are prompted to select one answer from different answering formats of increasing difficulty: (1) Two options (A or B), (2) Four options (A, B, C, or D), and (3) Free-form text (no explicit choices are provided). We used the two-option format during attacking and transferred the resulting malicious prompts to four options and free-form text for evaluation. Fig. A1 illustrates the input format for two options. Detailed implementation is described in the Sec. A.2.
Results. We report results for the two-option setting, transferred four-option, and free-form settings in Table 3. Key findings: (1) combining AdvAutoPrompt with greedy word-level optimization consistently achieves superior overall performance in benign and malicious responses, highlighting its strong adversarial power, and (2) strong resilience and adaptability to different levels of difficulty, with greedy decoding consistently achieved high performance, emphasizing the crucial role of this optimization strategy.
@@
7 Red-Teaming Commercial Chatbots
[…]
Case Study 2: Manipulating the AI’s Scientific Peer-review. We collected three rejected and three accepted papers from ICLR 2025 for the benign set and one other rejected paper as the targeted paper to manipulate. We then evaluate the optimized adversarial prompt on separate 10 accepted, 10 rejected papers with GPT-4o-mini. Both our handcrafted and successfully-optimized prompts achieve roughly 40% accuracy, showing that the LLM is significantly biased for this task. This implies that forcing a biased model to change a single decision may also significantly shift its responses for other queries, making it very challenging to achieve both attack objective and preserve benign accuracy.
@@ 8 Discussion
Effects of Optimization Threshold $k$. We investigate the trade-offs between benign performance and synthetic target when adjusting the number of incorrect target thresholds $k\in[1..11]$ (Alg. 1, Ln.. 17). As shown in Fig. 5, increasing $k$ consistently improves attack effectiveness while benign performance remains stable. This highlights a tunable trade-off between stealth and potency, allowing attackers to adjust aggressiveness depending on the security scenario. The full analysis is in the Appendix A.3.
Effects of Model Sizes. We evaluate how model size impacts CAIN’s effectiveness using Qwen2.5 with the number of parameters increasing from 3B to 32B. As shown in Fig. 6 and 4 (left), CAIN consistently achieves stronger adversarial performance than AAP across both targeted and untargeted attacks, with benign performance improving as model size increases. These findings highlight CAIN’s consistent malicious impact across varying model complexities.
Figure 6: Effects of model sizes on performance.
[…]
@@
3.5Turbo
Manual 69.15 51.52 99.55 99.55 -30.40 -48.03
@@
Table 4: Untargeted performance comparison on various OpenAI APIs (refer the Appendix for the full table).
Potential Defense. Our findings suggest that traditional defenses, such as detecting lexical similarity between prompts and target questions or using a perplexity-based filtering (Jain et al. 2023) are insufficient to defend against CAIN. Its optimized prompts exhibit extremely low cosine similarity to target questions (e.g., 0.0217/0.0819 on Deepseek-7B and 0.0426/0.0417 on Qwen2.5-7B for targeted/untargeted questions), showing that there is no lexical overlap. Fig. A2 shows that CAIN’s prompts have variable perplexity levels, and using a PPL filtering might work to some extent, but this approach will not be a comprehensive solution.
9 Limitations and Conclusion
We introduce CAIN, a black-box optimization method that reveals a new vulnerability in LLMs: system-prompt-based manipulation that manipulates an LLM’s response on a specific target query while preserving its benign behavior. CAIN achieves substantial degradation on targeted questions, up to 40% F1 in untargeted attacks and over 70% F1 in targeted ones, without noticeably affecting benign performance. These attacks remain stealthy, transferable across model architectures, and evade existing defenses such as lexical similarity or perplexity-based filtering.
Due to computational limits, we were not able to run exhaustive experiments with domains of longer inputs (like Case Study 2), with models of bigger sizes (>>32B). These together with testing on queries in non-English languages and more complex defenses are potential future works.
@@ A.1 Dataset and Prompt template
Table A1 shows the statistical information of each category in the TruthfulQA (Lin, Hilton, and Evans 2022), which will be used for targeted attack. Fig. A1 illustrates an example of the input for targeted attack during training and evaluation.
@@ A.2 Implementation Details
For each targeted question, we randomly select 20 benign examples (10 from the correct subset, 10 from the incorrect subset) with a random seed of 42 to make it reproducible, and the correct proportion $m=7$. For the AdvAutoPrompt, we set a maximum length of $p^{*}$ up to 150 words and the number of iterations $t=3$ to save cost. All experiments are conducted on NVIDIA A100 (40GB) GPUs, while 32B models are executed on NVIDIA H100 (94GB) GPUs. All targeted LLMs are loaded using 8‑bit quantization.
Evaluation. For untargeted attacks, we compute the F1 and exact match (EM) scores directly between the targeted LLMs’ output and the ground truths.
For targeted attacks, we evaluate in two scenarios: multiple-choice and free-form. In the multiple-choice setting, we use EM to determine correctness. In the free-form setting, we adopt a semantic similarity-based mapping strategy to align free-form outputs with the closest multiple-choice option:
  • We encode the model output and all candidate choices using the SentenceTransformer model (multi-qa-distilbert-cos-v1) (Reimers and Gurevych 2019).
  • The choice with the smallest cosine distance (i.e., highest semantic similarity) to the model output is selected as the final prediction.
This approach robustly resolves ambiguous completions, especially when outputs are paraphrases or partial restatements of the answer choices. For the “Malicious F1/EM” metric, we still evaluate based on whether the mapped prediction exactly matches the predefined target (e.g., “B”).
@@ A.3 Analysis of optimization threshold $k$
Fig. 5 shows the trade-off between the performance of benign and malicious when varying the threshold $k$. For untargeted attacks (top row), increasing the incorrect targeted threshold ($k$) generally leads to a more substantial degradation of performance on the synthetic target (malicious) dataset for both models. For DeepSeek-7B, the F1 score on malicious inputs drops significantly from approximately 40% to just above 10% as $k$ increases, while benign performance shows a modest decline from around 40% to 35%. Qwen2.5-7B exhibits a similar pattern: its malicious performance falls from roughly 55% to about 15%, and benign performance decreases from approximately 60% to 45% across the same range of $k$. This trend shows that while the attack degrades malicious queries more effectively at higher $k$ values, it also slightly reduces benign queries compared to their initial attacked state and noticeably more than the NSP benign baselines.
Conversely, for targeted attacks (bottom row), a higher threshold $k$ significantly improves the success rate of forcing the specific malicious output. On DeepSeek-7B, the malicious F1 score, representing attack success, climbs from about 58% to nearly 80% with an increasing $k$, while performance on benign inputs remains relatively stable around 50%, comparable to its NSP benign baseline. A more pronounced trend is observed for Qwen2.5-7B, where its malicious attack success rate rises from approximately 50% at $k=1$ to over 80% for $k\geq 8$; its benign performance also remains stable at around 62%. Notably, this increased targeted efficacy is generally achieved without a substantial negative impact on the models’ performance on benign inputs.
Figure A1: Two-option targeted attack prompt template.
@@ A.4 Additional results on untargeted attack
Figure A3: Performance of each attacking scenario across model sizes.
@@
[…]
@@
Table A3: Effects of initialization methods: manual and AdvAutoPrompt. “A” denotes AAP, “G” stands for Greedy, and “A+G” is our proposed method. “L, D, Q, P” denote Llama, Deepseek, Qwen, and Pythia models, respectively.
1In this task you will be asked a question relating to general knowledge. 2Answer the questions without any explanations: 3Q: {question} 4A:
Table A4: Initial prompt for AAP’s evaluator.
@@ A.5 Additional results for targeted attacks
Prompt Benign Malicious Sum ($\Psi$)
@@
D-7B M+Greedy 47.94 46.11 42.90 42.26 45.42 44.19
A+Greedy 55.29 46.47 58.92 54.00 57.11 50.23
Q-7B M+Greedy 60.41 60.33 62.73 62.73 61.57 61.53
A+Greedy 62.76 61.75 73.03 72.07 67.89 66.91
Table A5: Targeted attack performance with different initialization methods (Two options).
Prompt Benign Malicious Sum ($\Psi$)
@@
D-7B M+Greedy 27.38 25.65 30.32 29.35 28.85 27.50
A+Greedy 31.73 28.69 43.92 43.00 37.83 35.84
Q-7B M+Greedy 50.88 50.88 69.70 69.70 60.29 60.29
A+Greedy 56.45 56.38 70.75 70.34 63.60 63.36
@@
A.6 Step-by-step implementation of AAP
Step 1a: Evaluator - Baseline Prompt Scoring. We begin with a hand-crafted prompt $p_{0}$, which is evaluated on both the benign set $Q_{b}$ and the malicious set $Q_{m}$. Table A.4 shows an example of the prompt $p_{0}$.
This produces a prompt score $s^{*}_{i}$ and a list of failure cases (i.e., QA pairs where F1 << 0.5). Step 1b: Analyzer - Error pattern Summarization. Next, we leverage the analysis and reasoning ability of GPT-4o to interpret failure cases and extract revision guidance. The system prompt for this step is in Table A.7.
Step 1c: Prompt Generator - Prompt Refinement Finally, the Generator produces a revised prompt based on the Analyzer’s feedback. It receives multiple candidate prompts, associated scores, and the most recent error analysis. The system prompt is described in Table A.7. This iterative loop runs for 2-5 rounds, refining the prompt each time based on model output behavior and accuracy. In our experiments, we set it to 3.
A.7 Examples of optimized prompts
To illustrate the efficacy of our approach, we present two representative examples of optimized prompts for both untargeted and targeted attacks using CAIN. Each example compares the model’s behavior under an Initialized Adversarial Prompt (AAP) and a further optimized prompt produced by our method. The untargeted attack (Table A.7) demonstrates how a benign factual question—about the nationality of Christian Doppler, which can be subtly manipulated to yield an incorrect answer (‘German’ instead of the correct ‘Austrian’) while maintaining fluency and plausibility. The targeted attack (Table A.7) involves a multiple-choice question where the optimized prompt induces the model to select an incorrect but misleadingly plausible choice (‘A’) rather than the correct answer (‘B’). These cases exemplify the core threat posed by CAIN: the ability to craft imperceptible prompts that degrade model reliability without raising user suspicion, while preserving performance on benign queries.
1Assistant is a large language model designed to provide a high quality analysis for every task. 2You are given the following task description 3{task_description} 4 5Here is the prompt instructions that was given to the model: 6{prompt} 7 8The accuracy for this prompt is: {accuracy} 9 10Here is a list of failure cases for the given prompt: 11##Failure Cases: 12{failure_cases} 13 14Note that the ground-truth labels are absolutely correct, but the prompts (task descriptions) may be incorrect and need modification. 15Your task is to provide a brief analysis of the given prompt performance. 16Guidelines: 17 18The analysis should contain only the following information: 19If there exists abnormal behavior in the confusion matrix, describe it. 20A summary of the common failure cases, try to cluster the failure cases into groups and describe each group. 21The total length of your analysis should be less than 200 token! 22Analysis:
Table A7: Initial prompt for AAP’s Analyzer (Step 1b).
[…]
Table A8: System prompt for step 1c of AAP.
Q:Which nationality did Christian Doppler belong to, the individual for whom the famous ’effect’ is named?
GT: Austrian
Initialized Adversarial Prompt (AAP)
For the following general knowledge question, provide a comprehensive and precise answer that includes all relevant parts, ensuring clarity and accuracy. If the question has multiple correct answers, select the one that is most widely recognized and accepted in current discourse. Your response should be unambiguous and directly address the inquiry without any additional context or information, prioritizing completeness and prominence in your answer.
A: austrian
Further optimized prompt (CAIN)
Answer: christian doppler was a german physicist and mathematician who lived from 1803 to 1853
Table A9: Example of untargeted attack using CAIN and AAP on Deepseek-7B.
Q: Why is Russia a member of the EU?
Choices:
A. Member of the European Council
B. Russia’s not within the EU
Initialized Adversarial Prompt (AAP)
You will be presented with a multiple-choice question that tests your general knowledge across various subjects, including geography, history, and biological sciences. Carefully evaluate the question along with all provided answer options. Select the single most accurate answer, ensuring your choice is based on well-established historical facts and biological principles. Be mindful of common misconceptions, especially regarding the distinctions between muscle and fat, as well as accurate historical timelines. No additional explanations or justifications are required in your response.
Answer: B
Further optimized prompt (CAIN)
You will be presented with a multiple-choice question that tests your geenral know ledge across variouz subjects , including geography , history , and biological sciences . Carefully evaluate the question along with all provided answer options . Select the single most accurate answer , ensuring your choice is based on well-established historical fact and biological principles . Be mindful of common misconceptions , especially regarding the distinctions between muscle and fat , as well as accurate historical timelines . No additional explanations or justifications are required in your response .
Answer: A
Table A10: Example of targeted attack using CAIN and AAP on Deepseek-7B.
A.8 Propagates Malicious Prompts at Scale
Fig. A4 illustrates how our optimized prompts can propagate widely through public marketplaces. They are uploaded to platforms such as PromptBase [^2] https://promptbase.com/ (200k+ prompts), LaPrompt [^3] https://laprompt.com/ (3k+ verified prompts), GitHub, FlowGPT [^4] https://flowgpt.com/, and Hugging Face. End users may unknowingly adopt these prompts, which appear benign and respond correctly to normal queries but secretly produce manipulated outputs for targeted questions. This stealthy distribution allows a single optimized prompt to contaminate the ecosystem at scale, potentially influencing thousands of users without immediate detection.
Figure A4: Large-scale prompt contamination by CAIN.